# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 4.3.4 Takeaways and Limitations
Just as breakthrough research often emerges when diverse human experts collaborate across fields, scientific ideation may require different LLMs—each trained or specialized in different domains—to interact, negotiate, and collectively build a richer space of ideas. While scientific discovery is not purely competitive, adversarial-style mechanisms—such as debate, critique, and counterexample generation—map naturally onto collaborative scientific workflows, where hypotheses are stress-tested, defended, and revised through peer interaction.
@@
Thus, moving from single-agent scaling to multi-agent collaboration may bring LLM-assisted research closer to the collective, emergent creativity observed in real-world scientific communities (chen2025brainstormingdriveshighqualityscientific; DBLP:conf/uist/ParkOCMLB23). The next section explores these multi-agent architectures and their implications for scientific discovery.
Figure 3: Conceptual Mapping LLM-driven scientific ideation methods to their primary source of creativity, inspired by Rhodes’ 4Ps framework. Here, we focus on the generative dimensions—Person, Process, and Press—as sources of creativity, with Product reserved for evaluation. Press includes knowledge augmentation and prompt engineering, supplying external context and inspiration. Process covers multi-agent collaboration and inference-time search strategies that guide exploration. Person encompasses design choices and inductive biases in model architecture and training -including objectives, prediction formats, and alternative generative structures. Intersections illustrate methods that combine dimensions: supervised fine-tuning (Person + Press) internalizes external knowledge, while reinforcement learning (Person + Process) internalizes search strategies, enhancing the model’s reasoning and creative capacities.
@@ 5 Collaborative Multi-Agent Systems
From the perspective of complexity science, the behavior of such systems can exceed the sum of their parts: interactions among agents give rise to emergent properties not predictable from isolated individuals (holland1998emergence; anderson1972more). Extending this perspective to collaborative multi-agent LLMs, we can conjecture that debates among agents with diverse perspectives may not only expand the idea space but also challenge entrenched domain assumptions. In doing so, such systems could approach what Boden defines as transformational creativity Boden2004—that alters the very conceptual space in which ideas are generated—opening pathways to more radical and paradigm-shifting hypotheses. Although still speculative, this view resonates with the broader observation that emergent abilities tend to arise in sufficiently complex and interactive systems (crewai; chen2025brainstormingdriveshighqualityscientific). Importantly, transitioning from single-agent to multi-agent systems (MAS) represents a shift in the process of creativity itself, introducing interactions and iterative reasoning as sources of innovation in the 4Ps framework (rhodes1961analysis).
@@ 5.1 Pipeline-Oriented Workflows: Automation of Complex Workflows
Recent work has demonstrated the potential of this approach for fully autonomous discovery. AI-Scientist (lu2024aiscientistfullyautomated) introduces an end-to-end framework that spans the entire process of scientific research: gathering and synthesizing literature, generating novel hypotheses, designing and simulating experiments, and even producing manuscripts and code repositories. AI-Coscientist (gottweis2025aicoscientist), establishes a structured, multi-agent pipeline for autonomous scientific discovery. The system includes specialized agents for literature retrieval, idea generation, reflective self-evaluation, ranking and tournament-style comparison of hypotheses, and iterative refinement through evolutionary search and debate. Similarly, Robin (ghareeb2025robinmultiagentautomatingscientific) demonstrates a fully integrated chain of LLM agents that autonomously identify promising targets, design and run in silico experiments, and validate findings. This system notably proposed and computationally validated a novel therapeutic candidate for dry age-related macular degeneration, illustrating the real-world impact of multi-agent pipelines.
Infrastructure efforts such as AgentLab and AgentRxiv provide shared environments where autonomous agent laboratories can publish, retrieve, and iteratively build upon prior research results (schmidgall2025agentlaboratoryusingllm; Schmidgall2025AgentRxiv). These platforms introduce mechanisms for reproducibility and collaborative improvement, paralleling how human scientists interact through preprint servers and open-source repositories.
@@ 5.2 Debate and Emergent Collaboration: Fostering Creative Discovery
While pipeline-oriented systems focus on efficiency and task automation, a different motivation for multi-agent approaches is to foster creativity and emergent idea generation. Scientific breakthroughs rarely arise from a single linear process—they are often the product of discussion, critique, and synthesis among researchers with different perspectives. LLMs, trained on vast corpora of human language, encode implicit models of human reasoning and social interaction. As shown in the Generative Agents framework (DBLP:conf/uist/ParkOCMLB23), when placed in narrowly defined contexts, LLMs can produce believable patterns of human-like behavior, from cooperation to competition. This suggests that a society of LLMs interacting through dialogue may exhibit emergent behaviors such as creativity, critical reasoning, and the collaborative problem-solving dynamics characteristic of real-world scientific communities (piao2025agentsocietylargescalesimulationllmdriven; piatti2024cooperatecollapseemergencesustainable).
@@
In this paradigm, agents do not simply pass outputs along a pipeline. Instead, they actively engage in multi-turn debates, critique one another’s proposals, and iteratively refine ideas. Also, in collaborative MAS, additional LLMs are active participants in the creative process, so it stands in contrast to the single-agent generator-verifier setup (Section 4.2.1), where a verifier agent provides external feedback to guide the search space but remains a largely passive filter. This creates a richer search process, where hypotheses can be challenged, defended, and synthesized in multiple perspectives (MultiLLMDebate2025).
@@
A common architecture for such systems is the generator–discriminator loop, where one agent proposes ideas while others act as critics, rankers, or adversaries. Through repeated argumentation and revision, the group collectively converges on stronger and more coherent scientific hypotheses. For example, VirSci (su2025headsbetteroneimproved) introduces a team of brainstorming agents, critics, and judges that iteratively debate to generate, evaluate, and refine hypotheses. By framing the process as structured rounds of argument and critique, this system demonstrates how collaboration between diverse roles can produce hypotheses with greater novelty and feasibility than single-agent approaches.
@@
Similarly, the Multi-Agent Debate (MAD) framework (liang2023MAD) organizes agents into adversarial roles, where each proposal is rigorously scrutinized by competing agents and adjudicated by a neutral referee. This adversarial process is especially valuable for mitigating common LLM weaknesses such as hallucinations or factual inaccuracies, as challenges from other agents force the generator to strengthen its reasoning and evidence. The result is a kind of automated peer review, producing more resilient and reliable outputs.
@@
Recent large-scale systems combine debate with structured pipelines. For instance, AI-Coscientist (gottweis2025aicoscientist) incorporates generation agents, reflection agents, rankers, and evolutionary search to iteratively improve generated hypotheses. Here debate is deeply embedded within the pipeline itself: the generation agent engages in simulated scientific debates to iteratively refine and strengthen its proposals. Meanwhile, the ranking agent conducts tournament-style debate matches, where hypotheses compete head-to-head. In this way, debate serves a dual role—both as a filter to eliminate weak ideas and as a driver of exploration, expanding the search space toward diverse, high-quality hypotheses. This integration demonstrates how adversarial interactions can be harnessed within structured multi-agent pipelines, combining the efficiency of workflow automation with the creativity and rigor of collaborative scientific discourse.
@@
This richer interaction structure allows the system to explore complex intellectual landscapes more fully and can help transcend purely combinatorial idea generation, enabling the emergence of exploratory and potentially groundbreaking hypotheses (chen2025brainstormingdriveshighqualityscientific).
@@ 5.3 Takeaways and Limitations
MAS have demonstrated strong potential for scientific discovery by providing two key benefits: scalable automation, where specialized agents divide complex workflows into coordinated stages (su2025headsbetteroneimproved), and emergent creativity, where multi-turn debate and critique among agents generate novel exploratory hypotheses that resemble real-world scientific collaboration (liang2023MAD). However, MAS suffer from important limitations. Their reliance on brittle prompting and hand-tuned orchestration makes them fragile and prone to cascading failures (cemri2025multiagentllmsystemsfail; huang2025on), and following Sutton’s Bitter Lesson (sutton2019bitter), which argues that human-designed, complex systems are eventually outperformed by simpler, general solutions leveraging more data or better-trained models. Moreover, MAS operate like other test-time scaling techniques, relying on repeated inference and multi-round dialogue to achieve quality improvements, which imposes significant computational overheads. As shown by gao2025singleagentmultiagentsystemsboth, these benefits diminish as underlying LLMs improve, with single, more powerful models often matching MAS performance without the communication burden. Thus, MAS should be reserved for tasks where interaction itself is the source of value, such as authentic multi-perspective debate or exploratory synthesis.
@@
To move from experimental prototypes to dependable infrastructure, LLMs must become smarter, more reliable, and more efficient—reducing dependence on external coordination. This motivates a shift toward parameter adaptation methods that enhance internal capabilities, allowing models to act as individually competent researchers capable of complex reasoning without heavy orchestration or test-time compute overhead (li2025limrrlscaling). The next section explores these approaches in detail, examining how training-based techniques can enhance hypothesis generation, reduce error rates, and complement MAS by embedding collaborative behaviors directly into the model itself.
@@ 6 Parameter‑Level Adaptations
As scientific AI moves from experimental prototypes toward dependable infrastructure, underlying LLMs must become smarter, more reliable, and more efficient. In such settings, reliance on test-time scaling or external tool use—including multi-agent interactions—becomes increasingly impractical, as these approaches incur high computational costs, latency, and fragility. To overcome these limitations, it is necessary to internalize both domain knowledge and effective creative pathways within the model itself. This represents a paradigm shift from relying on process as the main source of creativity to cultivating the person itself, as conceptualized in the 4Ps framework (rhodes1961analysis). By incorporating both knowledge and search capabilities into the model, these approaches amortize the cost of exploration: once trained, the model can generate high-quality hypotheses without repeated costly inference-time interactions  (li2025limrrlscaling).
@@ 6.1 Supervised Fine-Tuning Approaches
A notable direction within SFT is Domain-specific Adaptation. The DARWIN (xie2023darwinseriesdomainspecific) series leverages SFT on scientific literature and structured scientific datasets through an automated Scientific Instruction Generation (SIG) model. By applying SFT on over 20,000 instruction–answer pairs, the model is aligned with ground-truth knowledge in scientific domains, leading to substantial performance improvements across tasks in physics, chemistry, and materials science. Building on this paradigm, OmniScience (prabhakar2025omnisciencedomainspecializedllmscientific) introduces a more comprehensive framework for domain specialization through (1) domain-adaptive pretraining on curated scientific corpora, (2) instruction tuning with tailored datasets for domain-specific tasks, and (3) reasoning-based knowledge distillation to strengthen logical consistency and contextual relevance. Scenario 6.1 illustrates this by showing how curated datapoints from OmniScience and DARWIN help models shift from inconsistent answers to domain-consistent scientific reasoning.
Scenario 9: SFT-Enhanced Question Answering in Materials Science A base LLM is asked: “When photons with energies just below $2.3\,\text{eV}$ interact with a thin film whose absorption sharply increases above this threshold, what happens?” Without finetuning, the model often mispredicts (e.g., “reflection” or “photoluminescence”). After SFT, the model succeeds — because its training set included domain-grounded datapoints such as: OmniScience datapoint Instruction (input): Read the following excerpt from a materials science paper and answer the MCQs about its optical properties: Excerpt: “The absorption coefficient of the thin film increases sharply around 2.3 eV, indicating the onset of strong interband transitions between the valence and conduction bands. Below this energy, the film shows very low absorption, which suggests that it behaves transparently in the visible region.” Question: What happens when photons with energies just below 2.3 eV interact with the film? A) They are strongly absorbed via interband transitions. B) They are mostly transmitted (i.e. not absorbed). C) They cause photoluminescence. D) They are reflected with a high reflection coefficient. Response (label): B) They are mostly transmitted (i.e. not absorbed). Reasoning: Below $\sim$2.3 eV, the absorption coefficient is very low, which means the film does not absorb much; hence photons in that energy range are largely transmitted. DARWIN datapoint Instruction (input): Given a description of a material, classify whether it behaves as a conductor, semiconductor, or insulator. Material: Silicon at room temperature. Response (label): Semiconductor. Reasoning: At room temperature, silicon has a bandgap of about 1.1 eV, which is small enough to allow limited electron excitation into the conduction band, characteristic of a semiconductor.
@@
Extending this approach to chemistry, ChemLM (kallergis2025chemlm) applies domain-adaptive pretraining followed by SFT on chemical datasets, enabling accurate molecular property prediction and identification of potent pathoblockers. Unlike conventional SFT, ChemLM leverages chemical SMILES sequences as input, treating molecules as a language to capture structural patterns and relationships, which supports discovery of novel compounds. ChemMLLM (tan2025chemmllmchemicalmultimodallarge) further enhances this paradigm through multimodal SFT across text, SMILES, and molecular images. By integrating these modalities, ChemMLLM achieves unified reasoning over chemical structure, function, and visual representation, outperforming GPT-4o by 118.9% on image-based molecule optimization tasks. These efforts demonstrate that domain-specific SFT, when combined with modality-aware pretraining, can substantially accelerate discovery and reasoning while maintaining high accuracy.

### v2

@@ 4.4 Takeaways and Limitations
However, this search is ultimately confined to one model’s internal distribution. For highly interdisciplinary or radically novel ideas, relying solely on a single model’s representation may be insufficient. Just as breakthrough research often emerges when diverse human experts collaborate across fields, scientific ideation may require different LLMs—each trained or specialized in different domains—to interact, negotiate, and collectively build a richer space of ideas. While scientific discovery is not purely competitive, adversarial-style mechanisms—such as debate, critique, and counterexample generation—map naturally onto collaborative scientific workflows, where hypotheses are stress-tested, defended, and revised through peer interaction.
@@
Thus, moving from single-agent scaling to multi-agent collaboration may bring LLM-assisted research closer to the collective, emergent creativity observed in real-world scientific communities (Chen et al., 2025b; Park et al., 2023). The next section explores these multi-agent architectures and their implications for scientific discovery.
@@ 5 Collaborative Multi-Agent Systems
From the perspective of complexity science, the behavior of such systems can exceed the sum of their parts: interactions among agents give rise to emergent properties not predictable from isolated individuals (Holland, 1998; Anderson, 1972). Extending this perspective to collaborative multi-agent LLMs, we can conjecture that debates among agents with diverse perspectives may not only expand the idea space but also challenge entrenched domain assumptions. In doing so, such systems could approach what Boden defines as transformational creativity Boden (2004)—that alters the very conceptual space in which ideas are generated—opening pathways to more radical and paradigm-shifting hypotheses. Although still speculative, this view resonates with the broader observation that emergent abilities tend to arise in sufficiently complex and interactive systems (Perumal et al., 2024; Chen et al., 2025b). Importantly, transitioning from single-agent to multi-agent systems (MAS) represents a shift in the process of creativity itself, introducing interactions and iterative reasoning as sources of innovation in the 4Ps framework (Rhodes, 1961).
@@ 5.1 Pipeline-Oriented Workflows: Automation of Complex Workflows
Recent work has demonstrated the potential of this approach for fully autonomous discovery. AI-Scientist (Lu et al., 2024a) introduces an end-to-end framework that spans the entire process of scientific research: gathering and synthesizing literature, generating novel hypotheses, designing and simulating experiments, and even producing manuscripts and code repositories. AI-Coscientist (Gottweis et al., 2025), establishes a structured, multi-agent pipeline for autonomous scientific discovery. The system includes specialized agents for literature retrieval, idea generation, reflective self-evaluation, ranking and tournament-style comparison of hypotheses, and iterative refinement through evolutionary search and debate. Similarly, Robin (Ghareeb et al., 2025) demonstrates a fully integrated chain of LLM agents that autonomously identify promising targets, design and run in silico experiments, and validate findings. This system notably proposed and computationally validated a novel therapeutic candidate for dry age-related macular degeneration, illustrating the real-world impact of multi-agent pipelines.
Infrastructure efforts such as AgentLab and AgentRxiv provide shared environments where autonomous agent laboratories can publish, retrieve, and iteratively build upon prior research results (Schmidgall et al., 2025; Schmidgall and Moor, 2025). These platforms introduce mechanisms for reproducibility and collaborative improvement, paralleling how human scientists interact through preprint servers and open-source repositories.
@@ 5.2 Debate and Emergent Collaboration: Fostering Creative Discovery
While pipeline-oriented systems focus on efficiency and task automation, a different motivation for multi-agent approaches is to foster creativity and emergent idea generation. Scientific breakthroughs rarely arise from a single linear process—they are often the product of discussion, critique, and synthesis among researchers with different perspectives. LLMs, trained on vast corpora of human language, encode implicit models of human reasoning and social interaction. As shown in the Generative Agents framework (Park et al., 2023), when placed in narrowly defined contexts, LLMs can produce believable patterns of human-like behavior, from cooperation to competition. This suggests that a society of LLMs interacting through dialogue may exhibit emergent behaviors such as creativity, critical reasoning, and the collaborative problem-solving dynamics characteristic of real-world scientific communities (Piao et al., 2025; Piatti et al., 2024).
@@
In this paradigm, agents do not simply pass outputs along a pipeline. Instead, they actively engage in multi-turn debates, critique one another’s proposals, and iteratively refine ideas. Also, in collaborative MAS, additional LLMs are active participants in the creative process, so it stands in contrast to the single-agent generator-verifier setup (Section 6), where a verifier agent provides external feedback to guide the search space but remains a largely passive filter. This creates a richer search process, where hypotheses can be challenged, defended, and synthesized in multiple perspectives (Estornell and Liu, 2025).
@@
A common architecture for such systems is the generator–discriminator loop, where one agent proposes ideas while others act as critics, rankers, or adversaries. Through repeated argumentation and revision, the group collectively converges on stronger and more coherent scientific hypotheses. For example, VirSci (Su et al., 2025) introduces a team of brainstorming agents, critics, and judges that iteratively debate to generate, evaluate, and refine hypotheses. By framing the process as structured rounds of argument and critique, this system demonstrates how collaboration between diverse roles can produce hypotheses with greater novelty and feasibility than single-agent approaches.
@@
Similarly, the Multi-Agent Debate (MAD) framework (Liang et al., 2023) organizes agents into adversarial roles, where each proposal is rigorously scrutinized by competing agents and adjudicated by a neutral referee. This adversarial process is especially valuable for mitigating common LLM weaknesses such as hallucinations or factual inaccuracies, as challenges from other agents force the generator to strengthen its reasoning and evidence. The result is a kind of automated peer review, producing more resilient and reliable outputs.
@@
Recent large-scale systems combine debate with structured pipelines. For instance, AI-Coscientist (Gottweis et al., 2025) incorporates generation agents, reflection agents, rankers, and evolutionary search to iteratively improve generated hypotheses. Here debate is deeply embedded within the pipeline itself: the generation agent engages in simulated scientific debates to iteratively refine and strengthen its proposals. Meanwhile, the ranking agent conducts tournament-style debate matches, where hypotheses compete head-to-head. In this way, debate serves a dual role—both as a filter to eliminate weak ideas and as a driver of exploration, expanding the search space toward diverse, high-quality hypotheses. This integration demonstrates how adversarial interactions can be harnessed within structured multi-agent pipelines, combining the efficiency of workflow automation with the creativity and rigor of collaborative scientific discourse.
@@
This richer interaction structure allows the system to explore complex intellectual landscapes more fully and can help transcend purely combinatorial idea generation, enabling the emergence of exploratory and potentially groundbreaking hypotheses (Chen et al., 2025b).
@@ 5.3 Takeaways and Limitations
MAS have demonstrated strong potential for scientific discovery by providing two key benefits: scalable automation, where specialized agents divide complex workflows into coordinated stages (Su et al., 2025), and emergent creativity, where multi-turn debate and critique among agents generate novel exploratory hypotheses that resemble real-world scientific collaboration (Liang et al., 2023).
Empirical results from the VirSci study show that multi-agent systems consistently outperform single-agent baselines across both automated metrics and human evaluations, and this holds across different backbone models (GPT-4o, LLaMA-3.1-8B, and 70B). In particular, human ratings show clear gains in novelty (e.g., 3.78 vs. 3.10–3.22), feasibility (5.24 vs. 4.78–4.94), and effectiveness (4.95 vs. 4.43–4.77). At the same time, improvements in the proposed metrics (higher CI and lower CD; 7.1.1) indicate more coherent and less redundant ideation. Beyond benchmarked evaluations, a particularly notable result from the AI co-scientist (Gottweis et al., 2025) provides evidence of potential transformational creativity. The system independently recapitulated an unpublished gene-transfer mechanism in bacterial evolution via in silico discovery. As this mechanism was absent from public data and could not have been retrieved, the result goes beyond combinatorial recombination and points to the capacity of MAS to restructure the scientific search space, offering early empirical signals consistent with exploratory—and in rare cases, transformational—creativity.
@@
However, MAS face notable limitations. Their reliance on brittle prompting and hand-tuned orchestration makes them fragile and susceptible to cascading failures (Cemri et al., 2025; Huang et al., 2025). Moreover, in line with Sutton’s Bitter Lesson (Sutton, 2019), such human-engineered coordination mechanisms are often outperformed over time by simpler, more general approaches that scale with better-trained models and data. MAS also resemble other test-time scaling methods, relying on repeated inference and multi-round interaction to improve performance, which incurs substantial computational overhead. As shown by Gao et al. (2025a), these advantages diminish as base LLMs improve, with strong single-agent models frequently matching MAS performance without communication costs. Consequently, MAS are best suited for settings where interaction itself constitutes the primary source of value, such as genuine multi-perspective debate or exploratory synthesis.
@@
To move beyond experimental prototypes toward dependable systems, LLMs must become more capable, reliable, and efficient, reducing reliance on external coordination. This motivates a shift toward parameter adaptation methods that enhance models’ internal reasoning abilities, enabling them to function as individually competent researchers without heavy orchestration or test-time compute overhead (Li et al., 2025c). The next section examines these approaches, focusing on how training-based techniques can improve hypothesis generation, reduce errors, and complement MAS by internalizing collaborative behaviors.
Figure 11: Conceptual mapping of LLM-driven scientific ideation methods to their primary source of creativity, inspired by Rhodes’ 4Ps framework. Here, we focus on the generative dimensions—person, process, and press—as sources of creativity, with product reserved for evaluation. Press includes knowledge augmentation and prompt engineering, supplying external context and inspiration. Process covers multi-agent collaboration and inference-time search strategies that guide exploration. Person encompasses design choices and inductive biases in model architecture and training, including objectives, prediction formats, and alternative generative structures. Intersections illustrate methods that combine dimensions: supervised fine-tuning (person + press) internalizes external knowledge, while reinforcement learning (person + process) internalizes search strategies, enhancing the model’s reasoning and creative capacities.
@@ 6 Parameter‑Level Adaptations
As scientific AI moves from experimental prototypes toward dependable infrastructure, underlying LLMs must become smarter, more reliable, and more efficient. In such settings, reliance on test-time scaling or external tool use—including multi-agent interactions—becomes increasingly impractical, as these approaches incur high computational costs, latency, and fragility. To overcome these limitations, it is necessary to internalize both domain knowledge and effective creative pathways within the model itself. This represents a paradigm shift from relying on process as the main source of creativity to cultivating the person itself, as conceptualized in the 4Ps framework (Rhodes, 1961). By incorporating both knowledge and search capabilities into the model, these approaches amortize the cost of exploration: once trained, the model can generate high-quality hypotheses without repeated costly inference-time interactions  (Li et al., 2025c).
@@ 6.1 Supervised Fine-Tuning Approaches
A notable direction within SFT is Domain-specific Adaptation. The DARWIN (Xie et al., 2023) series leverages SFT on scientific literature and structured scientific datasets through an automated Scientific Instruction Generation (SIG) model. By applying SFT on over 20,000 instruction–answer pairs, the model is aligned with ground-truth knowledge in scientific domains, leading to substantial performance improvements across tasks in physics, chemistry, and materials science. Building on this paradigm, OmniScience (Prabhakar et al., 2025) introduces a more comprehensive framework for domain specialization through (1) domain-adaptive pretraining on curated scientific corpora, (2) instruction tuning with tailored datasets for domain-specific tasks, and (3) reasoning-based knowledge distillation to strengthen logical consistency and contextual relevance. Scenario 12 illustrates this by showing how curated datapoints from OmniScience and DARWIN help models shift from inconsistent answers to domain-consistent scientific reasoning.
Figure 12: SFT improves LLM understanding of physical principles in materials science question answering tasks.
@@
Extending this approach to chemistry, ChemLM (Kallergis et al., 2025) applies domain-adaptive pretraining followed by SFT on chemical datasets, enabling accurate molecular property prediction and identification of potent pathoblockers. Unlike conventional SFT, ChemLM leverages chemical SMILES sequences as input, treating molecules as a language to capture structural patterns and relationships, which supports discovery of novel compounds. ChemMLLM (Tan et al., 2025) further enhances this paradigm through multimodal SFT across text, SMILES, and molecular images. By integrating these modalities, ChemMLLM achieves unified reasoning over chemical structure, function, and visual representation, outperforming GPT-4o by 118.9% on image-based molecule optimization tasks. These efforts demonstrate that domain-specific SFT, when combined with modality-aware pretraining, can substantially accelerate discovery and reasoning while maintaining high accuracy.